### v1

# GPT4Point: A Unified Framework for Point-Language Understanding and Generation

Abstract
Multimodal Large Language Models (MLLMs) have excelled in 2D image-text comprehension and image generation, but their understanding of the 3D world is notably deficient, limiting progress in 3D language understanding and generation. To solve this problem, we introduce GPT4Point, an innovative groundbreaking point-language multimodal model designed specifically for unified 3D object understanding and generation within the MLLM framework. GPT4Point as a powerful 3D MLLM seamlessly can execute a variety of point-text reference tasks such as point-cloud captioning and Q&A. Additionally, GPT4Point is equipped with advanced capabilities for controllable 3D generation, it can get high-quality results through a low-quality point-text feature maintaining the geometric shapes and colors. To support the expansive needs of 3D object-text pairs, we develop Pyramid-XL, a point-language dataset annotation engine. It constructs a large-scale database over 1M objects of varied text granularity levels from the Objaverse-XL dataset, essential for training GPT4Point. A comprehensive benchmark has been proposed to evaluate 3D point-language understanding capabilities. In extensive evaluations, GPT4Point has demonstrated superior performance in understanding and generation.
Figure 1: Task examples of GPT4Point. It performs accurate 3D recognition, detailed captioning, precise Q&A, and high-quality controllable 3D generation. Additionally, GPT4Point excels in 3D anomalous object description, accurately assessing abnormal shapes like the multi-face object and the 3D generation failure case. It is a crucial ability in the assessment of generated 3D objects.
1 Introduction
The recent Large Language Models (LLMs) [41, 42, 43, 59, 9, 57, 64, 49] have demonstrated remarkable advancements in the field of natural language processing. Inspired by their powerful capabilities, researchers have also explored Multimodal LLMs (MLLMs), via adapting LLMs into various modalities like images [31, 36], audio [24, 17, 5] and videos [70, 7]. The proliferation of extensive image-text pair [6, 51] has crucially enabled 2D MLLMs i.e., Vision Language Models (VLMs) to interpret images through textual representations. Concurrently, there is a growing trend in utilizing these multimodal models for guiding text-to-image generation [16, 55, 52, 71]. This represents a form of compression and reconstruction, exploring how to accurately recover and even edit the input image using controllable image generation models. However, despite the impressive capabilities of MLLMs in handling multiple modalities, they still face significant limitations in understanding and accurately interpreting the 3D world, a critical need for various important downstream applications like intelligent robotics and augmented reality.
Recent efforts to develop 3D MLLMs [23, 82] have notable limitations. Some  [23, 82] prioritize the overall scene and focus primarily on the spatial coordinates of objects, often neglecting the geometric details of individual objects. This can lead to a limited understanding of the 3D world. Meanwhile, these methods generally convert 2D image features into 3D representations [23], which leads to a substantial loss of geometric accuracy. 3D geometry information is important in understanding. As shown at the bottom of Fig. 1, the VLM fails to recognize the four-sided face object while our GPT4Point can figure out the anomalies. Concurrent works focusing on utilizing 3D features directly exhibit notable limitations. PointBind [21] exhibits a deficiency in training and demonstrates restricted text referencing abilities due to the limited dataset. On the other hand, PointLLM [66] necessitates the training of the corresponding Language Model (LLM) component and does not possess the capability to expand into text generation.
We present GPT4Point [^1] First author is the intern at Shanghai AI Laboratory., a novel unified framework for point-language understanding and generation. GPT4Point introduces the 3D object MLLM, which is a groundbreaking language model that fully utilizes point clouds to perform various point-text tasks as shown in Fig. 1. We utilize a Bert-based Point-QFormer for point-text feature alignment. Aligned features are separately input into the LLMs for text inference tasks and Diffusion for 3D object generation tasks. It is worth noting that, given a low-quality point cloud feature as a condition, GPT4Point can generate higher-quality results while maintaining the geometric shapes and colors by using point-text aligned features called controllable text-to-3D.
To tackle the scarcity of object point-language data [58], we leverage the Objaverse-XL dataset [12, 11] to develop an automated, effective data annotation engine Pyramid-XL. It employs Vision Language Models (VLMs) for generating text annotations. Pyramid-XL solves the problem that VLMs can not understand multi-view images directly. By synthesizing captions from multi-views obtained by the VLMs, the text annotation is stratified into three hierarchical levels, ranging from low to high, ultimately leading to precise annotations. Apart from the data engine, we establish an object point-text benchmark for assessing point multimodal model capabilities in recognition and text inference tasks, such as 3D object point cloud captioning, and Q&A. This benchmark also provides a critical standard for evaluating 3D object generation, while current assessments often rely on qualitative judgments from rendered images without a direct evaluation in 3D space [47]. Only relying on rendering images may lead to misunderstanding, for instance, in the bottom right of Fig. 1, a failure case produced by 3D generation (a bear has two bodies), makes 2D VLMs and even humans fail to recognize its anomaly but our model can identify with anomalies easily.
Our paper makes three major contributions:
  • We present the unified framework for point-language understanding and generation GPT4Point, including the 3D MLLM for point-text tasks and controlled 3D generation.
  • Introducing the automated point-language dataset annotation engine Pyramid-XL based on Objaverse-XL, currently encompassing 1M pairs of varying levels of coarseness and can be extended cost-effectively.
  • Establishing a novel object-level point cloud benchmark with comprehensive evaluation metrics for 3D point cloud language tasks. This benchmark thoroughly assesses models’ understanding capabilities and facilitates the evaluation of generated 3D objects.
Figure 2: The model architecture of GPT4Point for training. In Stage1, we employ a Bert-based [13] Point-Q-Former for point-text feature alignment through three point-text tasks. Then, in Stage2, an LLM is appended to train the model’s text inference capabilities. A Point Cloud Diffusion is attached separately to train controlled text-to-3D generation which keeps the geometry shape and colors.
2 Related Work
Multi-modal large language models (MLLMs). Large Language Models (LLMs) have demonstrated robust capabilities in language comprehension, reasoning, and generalization [41, 42, 43, 59, 9, 57, 64, 49]. Building upon this, Multimodal Large Language Models (MLLMs) extend these reasoning skills to additional modalities such as image [14, 76, 15, 18, 78, 80], audio [24, 17, 5], and video [33, 70, 7]. Typically, MLLMs align target features with textual features and then integrate them with LLMs for various text inference tasks. Some train the whole architecture from scratch [25, 46] and others [29, 3, 36, 31, 10] utilize pretrained LLMs. In the realm of 3D MLLMs, existing models either rely on 2D image information [23, 82] or simply align low-quality textual phrases with points [20, 66]. To solve these problems, we introduce a novel 3D MLLM designed for diverse point-text tasks. Our model, featuring a Point Q-Former based on Bert [13], aligns two domain features and integrates an LLM for text-based reasoning tasks, advancing the field of 3D multimodal understanding.
Language-driven 3D object understanding. 3D point cloud multimodal models encompass a broad spectrum, generally categorized into those focusing on the entire scene containing multiple objects and those focusing on individual objects. The former places more emphasis on the relative positions of objects in the scene rather than their geometric shapes; Here, we primarily focus on the latter. In a self-supervised way, powerful backbones like PointBert [72] for object points have been obtained [72, 44]. Then, point cloud language pretraining attempts to align the point cloud modality and the text modality. Some methods [75, 26] try to convert point clouds to depth images for alignment with text using CLIP [48]. Tri-modal approaches such as ULIP [68, 69, 21, 78] integrate point cloud, text, and image data. However, these methods all exclusively use 2D images, either explicitly or implicitly. Our work differs by directly aligning 3D point-text modalities, completely removing the dependency on image data.
Text-to-3D generation. Text-to-image generation models have experienced significant advancements recently [50, 73], yet text-to-3D models face challenges due to limited 3D data availability. Current approaches often rely on optimizing Neural Radiance Fields (NeRF) representation [39] with Score-Distillation-Sampling (SDS) loss [47]. While these optimization-based methods [47, 34, 8, 63] still fall short in robustness, speed, and generalization. Alternatively, Point-E [40] and Shap-E [28] employ feed-forward 3D generative models trained on large, undisclosed 3D datasets, offering better generalization and faster processing. However, these models often produce random, uncontrollable outputs with low-quality textures. To solve these limitations, we leverage point-text features to enhance the controllability of feed-forward models. This approach uses a low-quality point-text feature as a condition that allows for maintaining specific shapes and colors, thereby enabling the generation of higher-quality 3D objects.
Figure 3: Pyramid-XL: An automated point-text annotation engine. Directly inputting images into VLMs yields unsatisfactory results. We propose a progressive annotation approach with 3 levels of granularity, leveraging results from the previous level for precise outcomes.
3 Methods
This section provides an overview of our data text annotation engine and model architecture. In Sec. 3.1, we introduce Pyramid-XL, our point-language dataset annotation engine, discussing its design, function, and the progression from low-quality descriptions to ultimately precise and detailed ones. Then, in Sec. 3.2, we delve into GPT4Point’s architecture, explaining how to align point and text and demonstrating how LLM and point diffusion models contribute to unified understanding and generation.
3.1 Point-Language Dataset Annotation Engine
The public release of the large-scale Objaverse dataset [12] and its successor Objaverse-XL [11] includes 800K and 10M objects respectively, providing a vast amount of 3D object data. However, these objects lack corresponding text descriptions. We plan to use the rendered images of the objects as input and obtain textual descriptions through a trained Vision Language Model (VLM), however, we find that direct input of multi-view images into the VLM does not enable it to understand their 3D structure and give precise descriptions, as shown in the top right of Fig. 3. Hence, Pyramid-XL employs a hierarchical pipeline, evolving from initial low-quality descriptions to achieve ultimately precise and detailed results.
Pyramid-XL
Single-View Caption (Level 1): We use the primary VLM model BLIP-2 [31] to generate concise descriptions, approximately 10 words in length, from a single-view rendered image. Multi-View Caption (Level 2): This level synthesizes multiple Level 1 descriptions by GPT-4 [42] to create comprehensive multi-view captions which has approximately 30 words. VLM Instruction Caption and QA Pair (Level 3): Utilizing the view with the highest CLIP score, selected from textual descriptions, we engage the advanced VLM to produce detailed dense captions and a corresponding QA dataset.
In terms of scale, Pyramid-XL is employed to annotate over 1M objects with Level 1 captions, 660K objects with Level 2 captions (same as Cap3D [38]), and 70K objects with Dense Captions including QA data. To assess the impact of text granularity on training, we designate the 1M Level 1 captions as the pretrain dataset, while a smaller set of detailed Level 3 data is used for instruction tuning. This methodology mirrors practices in the vision field, where models are initially pretrained on large volumes of coarser data and subsequently finetuned on more detailed data from specialized domains. Detailed experimental results of this approach are presented in Sec. 5.3.
3.2 Model Architecture
GPT4Point consists of two stages as illustrated in Fig. 2. In Stage1, we focus on point-text alignment using the Point-QFormer, a Bert-based structure similar to the Q-Former in BLIP-2 [31]. This stage involves supervision through three tasks related to recognition and text reasoning. In Stage2, only the point cloud is input into the point encoder and Point-QFormer to obtain aligned features, which are then devided into two branches: the LLM Branch and the Diffusion Branch separately. These branches supervise text comprehension and object generation tasks, respectively.
Stage1: point-text feature alignment. Given a point cloud ${\bm{P}}\in\mathbb{R}^{N\times 6}$, where each point is represented by six dimensions (XYZ coordinates and RGB color values), the initial stage of training focuses on feature extraction. The point encoder $\mathcal{E}$ processes the point cloud to yield the point cloud feature token ${\bm{T}}_{1}^{p}=\mathcal{E}({\bm{P}})$. Concurrently, the input text goes through tokenization via the Point Q-Former’s text tokenizer, resulting in the text feature token ${\bm{T}}_{1}^{t}$. These tokens, ${\bm{T}}_{1}^{p}$ and ${\bm{T}}_{1}^{t}$, are then utilized as inputs for the Point Q-Former $\mathcal{F}_{Q}$, facilitating the fusion of point cloud and textual data. We jointly optimize three training objectives: Point-Text Contrast (PTC) and Point-Text Matching (PTM), both recognition tasks, along with Point Caption Generation (PTG), a text inference task designed for aligning point clouds with textual data. The formulas are as follows:
$\mathcal{L}_{1}=\mathcal{F}_{Q}\left({\bm{T}}_{1}^{p},{\bm{T}}_{1}^{t}\right)=\mathcal{F}_{Q}\left(\mathcal{E}\left({\bm{P}}\right),{\bm{T}}_{1}^{t}\right)$ (1)
Here, $\mathcal{L}_{1}$ represents the loss for three tasks, and we have set the weight ratios between them all to 1. In the final layer of $\mathcal{E}$, a fully connected layer maintains consistency between the dimensions of ${\bm{T}}_{1}^{p}$ and ${\bm{T}}_{1}^{t}$.
Stage2: point understing and generation. After the point-text feature alignment, we proceed with understanding and generation tasks. It’s important to note that here we only input the point cloud into the Point Encoder and Point Q-Former to obtain the aligned feature. For the understanding task, a Large Language Model (LLM) is integrated with the Point Q-Former. The semantically integrated point cloud features are represented as ${\bm{T}}_{2}^{P}=\mathcal{F}_{Q}\left({\bm{T}}_{1}^{p}\right)=\mathcal{F}_{Q}\left(\mathcal{E}\left({\bm{P}}\right)\right)$. The textual feature tokens ${\bm{T}}_{2}^{t}$ are obtained from the LLM’s own tokenizer. The objective function is defined as follows:
[…]
$\mathcal{F}_{Q}$ indicates Point Q-former including a fully connected layer in its last layer to ensure consistency between the dimensions of ${\bm{T}}_{2}^{p}$ and ${\bm{T}}_{2}^{t}$. $\mathcal{L}_{2}$ represents the loss function from the Point Caption task alone.
For 3D object generation, we utilize the features obtained from low-quality point clouds via the Point Q-Former as conditions inputted into the text-to-3D framework. This process results in the generation of refined 3D objects that maintain consistency in shape and color with the original point cloud. A notable distinction from the LLM branch is that we have not only frozen point cloud diffusion but also frozen Point Q-Former. As shown in Fig. 2, we employ a single fully-connected layer to project the aligned features into the CLIP token embedding space, referred to as ${\bm{T}}_{3}^{p}$, and then concatenate these with the original text embeddings ${\bm{T}}_{3}^{t}$ using the CLIP tokenizer. The output from the CLIP text encoder, enriched with information from the original point cloud, is instrumental in enabling effective text-to-3D generation. The final output is achieved using Point-E. This framework is inspired by BLIP-Diffusion [30] techniques used in subject-driven 2D generation. However, the key distinction here from BLIP-Diffusion lies in the way we concatenate the Clip text token and Q-Former feature. This difference may also stem from variations in the data volumes between 2D and 3D, which will be thoroughly examined in the appendix.
Figure 4: Examples of text inference using the GPT4Point with ViT-g and OPT6.7B after Instruct Finetuning. The table showcases its proficiency with point cloud input, excelling in tasks like detailed caption generation and point cloud-based question answering. This underscores our model’s profound grasp of point cloud geometry and color, translating them into meaningful semantics.
4 Benchmarks and Evaluation
Evaluating the performance of multimodal models presents significant challenges due to the lack of mature metrics for assessing the quality of generated texts. For 3D objects, benchmarks primarily rely on human judgment or GPT-based assessments [66]. There are two key issues to consider in this context. Firstly, the evaluation process involves a certain degree of subjectivity. Identical results might receive varying scores, leading to an element of randomness. Secondly, each evaluation incurs time and monetary costs. In this section, we present the evaluation benchmark we have proposed, which is primarily designed to be objective, ensuring repeatability and verifiability. Sec. 4.1 outlines the composition of our test set. Sec. 4.2 addresses the evaluation of recognition capabilities, while Sec. 4.2 provides a detailed assessment of text inference abilities.
4.1 Composition of Test Set
We leverage the Objaverse dataset [12], aligning it with LVIS categories [22], to create Objaverse-LVIS validation and test sets. In Objaverse-LVIS, we exclude scenes with complex settings, such as indoor houses or outdoor parks, focusing more on scenarios with single objects or combinations of multiple objects. We construct validation and test sets, each containing 1K objects. Compared to the PointLLM [66], which uses only 200 unfiltered objects as a test set, our larger set of 1K objects better measures the model’s generalization capabilities. For textual descriptions, we initially use Pyramid-XL to get initial annotations, followed by multiple rounds of expert manual revisions, ensuring comprehensive and accurate descriptions.
4.2 3D Object Recognition
3D object recognition represents the classification capabilities of 3D multimodal models and the ability to match point cloud features with textual features. Objective measures, like accuracy, are typically used for evaluation.
Zero-shot point classification. Zero-shot point classification is considered a classic task in this domain. The widely used ModelNet40 dataset [65], which includes 2,468 objects across 40 categories, serves as a benchmark to evaluate a model’s classification capabilities. In the multimodal context, the typical approach involves using the text ’a 3D model of [name]’ as input to match with the point cloud modal features. The accuracy metric ACC@1, indicating the precision of top-1 rankings, best reflects the model’s ability to accurately match object categories.
3D point-text retrieval. In 3D Point-Text Retrieval, we initially select 128 candidates based on point-text feature similarity and then re-rank these candidates using matching scores. Unlike classification tasks where the text usually involves simple category names, here the text can be more complex descriptions. The evaluation metrics used are similar to those in image-text retrieval. We employ R1, R5, and R10 metrics to measure the accuracy of the top 1, 5, and 10 results in correctly matching points to text and vice versa.
4.3 3D Object Text Inference
3D object text inference deeply represents the understanding capabilities regarding objects, including 3D object point cloud captioning and 3D point cloud question answering.
3D point cloud captioning. This task primarily evaluates the model’s ability to provide an overall summary of a 3D object. The captions in the Objaverse-XL-LVIS caption test set are mostly within 30 words and accurately describe the object’s geometry, color, and state. And we predominantly employ common image description metrics, such as BLEU1, BLEU4, METEOR, ROGUE-L, and CIDEr [45, 35, 4] for evaluation.
3D point cloud question answering. In addition to point cloud captioning, 3D point cloud question answering explores object details through multiple rounds of dialogue. For instance, we can further explore the color or shape of specific parts of an object or even infer its simple usage. The curated Objaverse-XL-LVIS short QA 1K test set features concise, straightforward questions and answers, allowing us to conveniently calculate answer accuracy. Besides accuracy, we also use metrics from captioning to evaluate model performance. It is important to note that, for a fair comparison, we solely utilize zero-shot learning, meaning no fine-tuning is conducted on this kind of short QA dataset.
Table 1: Point-Text Retrieval on the Objaversexl-LVIS test dataset and zero-shot 3D classification on ModelNet40. Please note that denotes Generative 3D object classification, which refers to the process of classifying 3D objects based on the generation of captions.
Model Input Data Type ObjaverseXL-LVIS Retrieval (1K test set)
ModelNet40[65]
Point $\rightarrow$ Text Text $\rightarrow$ Point
Accuracy
R@1 R@5 R@10 R@1 R@5 R@10
Acc@1
  Image-Text Modal
BLIP-2
Single-View Image (Mesh with Color)
17.56 41.16 52.82 16.72 40.2 52.56
35.62
InstructBLIP 20.4 43.1 55.3 13.7 32.5 42.7
31.48
  Point-Text Modal
PointLLM(Vicuna-7B) Point Cloud (+Color) - - - - - -
41.33
GPT4Point 32.2 64 81.3 89.7 98.1 98.9 43.90
Model #Trainable Params ObjaverseXL-LVIS Caption (1K test set) ObjaXL-LVIS QA (1K)
BLEU1 BLEU4 METEOR ROUGH-L CIDEr Acc
BLEU1
ROUGH
  Image-Text Modal
BLIP-2 (OPT2.7B) 188M 22.2 3.0 10.3 28.2 32.3 13.4
14.2
16.8
BLIP-2 (OPT6.7B) 188M 24.9 4.1 11.5 30.0 44.2 15.4
15.1
18.3
InstructBLIP(Vicuna-1B) 202M 25.5 4.3 11.6 30.7 47.2 15.9
16.2
20.1
Qwen-VL(Qwen-7B) 7.2B 27.1 4.9 13.1 31.3 63.8 18.2
19.5
24.4
  Point-Text Modal
PointLLM (Vicuna-13B) 13.3B 26.2 4.9 11.9 31.3 50.9 23.4
22.3
26.2
GPT4Point (OPT2.7B) 110M 28.9 6.0 13.2 33.9 68.4 22.1
23.4
25.3
GPT4Point (OPT6.7B) 110M 31.5 7.2 13.8 35.4 78.7 27.1
26.2
30.4
GPT4Point (FLANT5XL) 110M 32.2 7.2 14.2 35.5 78.0 27.6 26.3 31.3
Table 2: 3D Object Point Caption and Question Answer (QA) on the Objaversexl-LVIS 1K test dataset. For the BLIP series, only fine-tuning of the Q-Former structure is required, whereas models like PointLLM need fine-tuning of the large language model.
Table 1: Point-Text Retrieval on the Objaversexl-LVIS test dataset and zero-shot 3D classification on ModelNet40. Please note that denotes Generative 3D object classification, which refers to the process of classifying 3D objects based on the generation of captions.